OpenStudy (anonymous):

sara is selling cookies and brownies as a fundraiser for the school band. she charges 2.25 for two cookies and one brownie. she charges 5.75 for four cookies and three brownies. how much does sara charge per cookie? how much per brownie? a. 0.50 per cookie; 1.25 per brownie b. 0.75 per cookie; 1.50 per brownie c. 0.25 per cookie; 1.00 per brownie d. 1.00 per cookie; 1.00 per brownie

OpenStudy (anonymous):

cookies = x brownies = y 2x + y = 2.25 y = 2.25 - 2x 4x +3y = 5.75 Solve the eq

OpenStudy (anonymous):

its $0.50 per cookie and $1.25 per brownie